OpenStudy (anonymous):

Find the solution of the differential equation y'=(5t+1)^2 satisfying the initial condition y(0)=4

OpenStudy (ddcamp):

Start by taking the integral of y': \[\int\limits (5t+1)^2dt\] Doing a u-sub, u = 5t + 1 du = 5 dt \[ \int\limits_{}{} u^2 (5)du\\ 5 \int\limits_{}{} u^2du \\ \frac{5}{3}u^3+C\] So, if y= (5/3)(5t+1)+C, and y(0) = 4, what must C be?

OpenStudy (anonymous):

Ah, beat me to it.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

idk honestly how would you get that?

OpenStudy (anonymous):

@bloopman can you help me?

OpenStudy (ddcamp):

\[\frac{ 5 }{ 3 }(5(0)+1)+C=4\\ \frac{5}{3}+C=4 \\ C = \frac{7}{3} \\. \\. \\. \\. \\ y = \frac{5}{3}(5x+1)+\frac{7}{3}\]

OpenStudy (anonymous):

is that it or their is more beyond this?

OpenStudy (ddcamp):

I don't know what notation your answer needs to be in, but I would recommend simplifying that final equation.

OpenStudy (anonymous):

the final answer has to be in "y=" form

OpenStudy (anonymous):

in the instructions y' means find the first derivative

OpenStudy (ddcamp):

If the answer needs to be in y= form, just simplify the answer we got after solving for C.

OpenStudy (anonymous):

25/3x+2/3?

OpenStudy (anonymous):

is that the simplified version?

zepdrix (zepdrix):

Woops Camp you missed a little something in your U-sub. \[\Large du=5dt \qquad\to\qquad \frac{1}{5}du=dt\]

zepdrix (zepdrix):

@DDCamp

OpenStudy (anonymous):

@zepdrix where do we go from there?

zepdrix (zepdrix):

So our substitution should actually give us:\[\Large \color{red}{\cancel{\color{black}{5 \int\limits u^2du}}}, \qquad\qquad\qquad \frac{1}{5}\int\limits u^2\;du\]Everything else looks ok though :o

zepdrix (zepdrix):

Confused about any of the steps?

OpenStudy (anonymous):

is it cool if you can go do the next steps ? :)

zepdrix (zepdrix):

Integrating, applying the power rule for integration:\[\Large \int\limits y'\quad=\quad \frac{1}{5}\int\limits u^2\;du\]gives us:\[\Large y\quad=\quad \frac{1}{5}\cdot \frac{1}{3}u^3+C\]

zepdrix (zepdrix):

\[\Large y(t)\quad=\quad \frac{1}{15}u^3+C\]Undoing our sub,\[\Large y(t)\quad=\quad \frac{1}{15}(5t+1)^3+C\]

zepdrix (zepdrix):

\[\Large y(\color{#F35633}{t})\quad=\quad \frac{1}{15}(5\color{#F35633}{t}+1)^3+C\]

zepdrix (zepdrix):

Using our initial data:\[\Large y(\color{#F35633}{0})\quad=\quad \frac{1}{15}(5\cdot\color{#F35633}{0}+1)^3+C\quad=\quad 4\]We can solve for C

zepdrix (zepdrix):

Mmm so what do we get for C?

OpenStudy (anonymous):

60/15

zepdrix (zepdrix):

When we plug 0 in for t, we should still get a 1 being cubed inside of those brackets.

OpenStudy (anonymous):

i meant 59/15

zepdrix (zepdrix):

Oh ok :)

zepdrix (zepdrix):

So we plug that in for our newly found C value, \[\Large y(t)\quad=\quad \frac{1}{15}(5t+1)^3+\color{orangered}{\frac{59}{15}}\]

zepdrix (zepdrix):

And we're done! yay

OpenStudy (anonymous):

thanks you soo much you came to the rescue haha

zepdrix (zepdrix):

:3